### v1

# Learning to detect continuous gravitational waves: an open data-analysis competition

Abstract
We report results of a public data-analysis challenge, hosted on the open data-science platform Kaggle, to detect simulated continuous gravitational-wave signals. These are weak signals from rapidly spinning neutron stars that remain undetected despite extensive searches. The competition dataset consisted of a population of CW signals using both simulated and real LIGO detector data matching the conditions of actual CW searches. The competition attracted more than 1,000 participants to develop realistic CW search algorithms. We describe the top 10 approaches and discuss their applicability as a pre-processing step compared to standard CW-search approaches. For the competition’s dataset, we find that top approaches can reduce the computing cost by 1 to 3 orders of magnitude at a $2\%$ dismissal probability. Additionally, the competition drove the development of new GPU-accelerated detection pipelines and extended CW-inspired statistics to other domains. We release the associated dataset, which constitutes the first open standardized benchmark for CW detection, to enable reproducible method comparisons and to encourage further developments toward the first detection of these elusive signals.
1 Introduction
Continuous gravitational waves (CWs) [1, 2, 3, 4] are long-duration gravitational-wave (GW) signals expected from rapidly-spinning non-axisymmetric neutron stars (NSs). Although currently undetected, their detection would, shed light on the extreme physics of NSs, allow us to understand their population in our galaxy, or even establish a new probe to direct and indirect detection of dark matter.
CW searches can be classified according to their assumptions on the expected source [5]. Targeted searches, for example, assume the CW emission of a pulsar fully matches its electromagnetic emission; this allows for very sensitive searches at the cost of not being robust to slight mismatches of the signal. On the other end of the spectrum, blind searches attempt to detect a CW signal from unknown sources; due to their parameter-space breadth, these kinds of searches cannot make use of matched-filtering and instead resolve to semicoherent methods to trade sensitivity for computing cost [6]. Incidentally, these approaches make blind searches more robust to un-modelled physics such as NS glitches [7] or spin-wandering [8, 9].
[…]
Machine learning applications [35] have demonstrated significant accelerations in search [36, 37, 38, 39] and parameter-estimation [40, 41, 42, 43] workflows targeting compact binary coalescences compared to standard approaches [44, 45, 46]. Due to the comparatively broader parameter space, applications to blind CW signals have reported limited success [47, 48, 49, 50, 51], with most successful applications focusing on accelerating specific computational steps [52] or increasing the robustness of a search to unmodelled physics [53, 54, 55].
In this work, we present the results of an open data-analysis competition to spark new approaches and algorithmic developments in CW searches [56]. This competition delivered a comparison of methods beyond what within-domain mock data challenges are able to prospect [57, 58, 59, 60] and allowed us to understand the complications and computational advantages achievable by off-the-shelf approaches from beyond the community. We release the dataset of the competition [61] to provide the first publicly-available benchmark targeted towards the detection of CW searches.
We also summarize the results of the first Kaggle competition aimed at detecting GW signals from compact binary coalescences [62] in A.
2 Leveraging open competitions for gravitational-wave research
The development of CW search methods is often constrained by the expertise and computational perspective available within the CW research community. The complexity of detecting a signal in noisy data, however, represents a challenge that may benefit, if properly supervised and tested, from a fresh perspective as provided by the broader data science community.
Open data-analysis competitions have demonstrated significant value in advancing scientific research across multiple domains, such as medical imaging [63, 64, 65], time-series forecasting [66], or particle physics [67], often producing solutions that outperform existing standards. These successes stem from competitions’ ability to attract diverse problem-solving approaches, encourage rapid iteration, and provide standardized benchmarks for method comparison [68].
The gravitational-wave community has already embraced crowd-sourcing approaches through platforms, such as Zooniverse [69], to understand transient noise in interferometric detectors [70, 71, 72], or Kaggle [73] to search for gravitational-wave signals from compact binary coalescences [62]. This approach can be understood as an evolution of the well-established mock data challenges conducted by the GW community [57, 58, 59, 60], where the aim has shifted from comparing existing pipelines to exploring new strategies altogether. The overall experience in GW citizen-science so far suggests non-experts, when given a proper training framework, can deliver solutions beneficial to the GW community [74, 39, 75].
The problem of detecting CW signals presents an ideal case for an open competition: First, the problem can be framed as a well-defined classification task [76, 77, 78], which provides a simple criterion to rank different strategies against each other. Second, synthetic data can be easily generated, either from first principles or through well-established software packages [79, 80, 81]. Third, since large-scale searches can be understood as the aggregation of many small-scale searches [29], competitors can tackle a simulation of the real problem using standard hardware. Finally, despite the existing experience in the field, large portions of the search-pipeline configuration space where novel improvements may be found remain yet to be explored.
Building on previous positive results [62], we released our competition on Kaggle [73], a well-established data-analysis platform powered by Google with proven track record in scientific applications and more than a million active users spread across vast domains of data science and engineering.
3 Dataset description
We seek a data-analysis strategy compatible with a blind CW search; that is, a pipeline which classifies narrow frequency bands of LIGO/Virgo/KAGRA [82, 83, 84, 85] data according to whether they contain a signal or not. We take the deep searches conducted using GCT [86] and Weave [87] as a gold standard for our competition. These pipelines use highly-efficient implementations of the semicoherent $\mathcal{F}$-statistic [20, 21, 23, 24, 22] to deliver the most sensitive all-sky search results across their explored parameter space. Their sensitivity is effectively bounded by the available computing resources, which amount to about 100 million CPU hours or 1 million GPU hours, depending on the specific implementation. By framing CW detection as a binary classification problem, we aim to discover whether alternative approaches could compete with or complement these well-established strategies.
3.1 Data format
We constructed a dataset using Short Fourier Transforms [88] (SFTs), a lossless data format used by actual blind CW searches [26, 23, 24, 53, 28, 25, 29]. This choice provides the participants with the same information available to CW data analysts and simplifies the transition of a successful solution into a production stage.
Given a set of $N$ time-domain data samples $x_{\alpha}[j],j=0,\dots,N-1$ starting at an epoch $t_{\alpha}$ sampled with a period $\Delta t$, we define an SFT following [88]:
$\tilde{x}_{\alpha}[k]=\Delta t\sum_{j=0}^{N-1}x_{\alpha}[j]e^{-\frac{-2i\pi jk}{N}}\,.$ (1)
Here the index $k=0,\dots,N/2$ labels the positive frequencies $f_{k}=k/T_{\textrm{SFT}}$, where $T_{\textrm{SFT}}=N\Delta t$ is the duration of an SFT. These frequencies physically correspond to $F_{0}+f_{k}$, where the fundamental frequency $F_{0}$ is always extracted through a homodyne filter before computing the SFT.
The competition dataset consisted of 8,000 test samples divided into signal and noise samples in equal parts. Additionally, 600 training samples were provided to the competitors as a quick entrypoint to the competition. The dataset was made available on the Kaggle platform, and adds up to $\sim\!0.5\,\mathrm{TB}$. Additionally, we supplied the tools to generate simulated CW data through the PyFstat package [81], which provides a Python interface to the LIGO/Virgo/KAGRA Algorithm Library [79, 80].
The dataset contains both noise-only samples and software-simulated CW signals, using both simulated Gaussian noise and data slices from the third observing run of the Advanced LIGO detectors [89, 90, 91]. The chosen frequency band ($0.2\text{\,}\mathrm{Hz}$ or $0.2\text{\,}\mathrm{Hz}$\times T_{\textrm{SFT}}=360$ frequency bins) is compatible with those analyzed by current blind searches. The sample duration, on the other hand, falls on the shorter end, as most searches analyse at least 6 months of data. This choice is due to the limited storage available on the Kaggle platform. Regardless, the duration is long enough to contain the main distinctive features of CW signals, namely a Doppler shift on the time-scale of months combined with a daily amplitude modulation.
3.2 Noise distribution
[…]
Each data sample in the competition’s dataset consists of a two sets of Short Fourier Transforms, representing the H1 (Hanford) and L1 (Livingston) LIGO detectors, spanning a total time of 4 months and a frequency band of $0.2\text{\,}\mathrm{Hz}$ with $T_{\textrm{SFT}}=$1800\text{\,}\mathrm{s}$\$. To better simulate the conditions of a real search, we introduce gaps in the data. As a result, SFTs are not necessarily contiguous (i. e. $t_{\alpha+1}-t_{\alpha}\geq T_{\textrm{SFT}}$).
[…]
To generate a data sample, we fix the total duration at $T_{\textrm{span}}=4\;\mathrm{month}$ and sample gap/science-mode durations alternatively. The first segment is sampled from the gap distribution and added to the starting time of O3 to avoid fixing the starting time. This process is repeated for both interferometers independently. By construction, the resulting data samples have a “duty cycle” (fraction of science-mode data) of
$\textrm{dc}=\frac{\tau_{\textrm{data}}}{\tau_{\textrm{data}}+\tau_{\textrm{gap}}}\simeq 0.78\,$, (2)
which is comparable with realistic observing runs [93, 94].
Figure 1: Distribution of observing-mode (solid line) and gap (dashed line) duration throughout the O3 observing run. Different colors refer to different detectors. Gray lines are an exponential fit as discussed in the text.
To generate a data sample containing real noise, we sample a subset of the full observing run amounting to $T_{\textrm{span}}=4\;\mathrm{month}$ following the distributions in Fig. 1. To avoid cross-correlation of strong instrumental artifacts with actual Advance LIGO data, we shift the frequency labels a random amount within $\pm 4\text{\,}\mathrm{Hz}$ and add random Gaussian noise with 0.1% of the data’s ASD.
3.3 Signal distribution
CW signals from isolated NSs are typically described in terms of the amplitude parameters $\mathcal{A}$, comprising the GW’s nominal amplitude $h_{0}$, the (cosine of the) inclination angle with respect to the line of sight $\cos\iota$, the polarisation angle $\psi$, and the initial phase $\phi_{0}$; and the phase-evolution parameters $\lambda$, which include the signal’s frequency $f_{0}$ and frequency derivative (spindown) $f_{1}$ at a fiducial reference time $t_{\mathrm{ref}}$, and the sky location $\hat{n}$. We take the reference time at the start of the O3 run and parameterise the sky position using equatorial coordinates (right ascension $\alpha$ and declination $\delta$).
We simulate a source population compatible with that assumed in blind searches, as summarized in Table 1. To generate a signal sample, we draw a set of parameters $(\mathcal{A},\lambda)$ from the population, we simulate noise as described in Sec. 3.2 for both the H1 and L1 LIGO detectors, and use MakeFakeData_v5 [79] to add a simulated signal into the noise. This software simulates a CW signal as observed by a ground-based interferometric detector, adds Gaussian noise with a given ASD, and returns a the corresponding SFTs with the specified $T_{\textrm{SFT}}$.
The sky position $\hat{n}$ is drawn uniformly across the celestial sphere; the distribution over the initial phase and orientation angles $(\cos\iota,\psi)$ follows that of an isotropically oriented-population. The spindown $f_{1}$ is drawn from a log-uniform distribution along $10^{-8},-10^{-12}]$ with 90% probability and $10^{-12},10^{-9}]$ with a 10% probability. This split accounts for possible spin-ups due to accretion, proper motion, or the evaporation of a boson cloud [20, 8, 95]
The frequency $f_{0}$ is sampled uniformly along $50,500]\,\mathrm{Hz}$. The maximum value of $500\text{\,}\mathrm{Hz}$ is chosen so that the CW signal does not drift more than one frequency bin within an SFT. The bandwidth of a CW signal depends on its frequency and sky location, as it is largely dominated by the Doppler effect due to the orbital motion of the detector around the Sun. For the parameter space considered in this challenge, the maximum bandwidth is about $0.05\text{\,}\mathrm{Hz}$. To complete a band of $0.2\text{\,}\mathrm{Hz}$, we attach noise data to both sides of the signal’s spectrum so it is contained completely within the data sample. Also, to avoid biasing the dataset, the proportion of data attached to each side of the signal’s frequency band is chosen randomly so that the signal is not centred.
Parameter Unit Symbol Distribution
[…]
$10^{-\mathrm{Uniform}(9,12)}\,(10\%)$
[…]
(sin) Declination $\sin{\delta}$ $\mathrm{Uniform}(-1,1)$
(cos) Inclination angle $\cos{\iota}$ $\textrm{Uniform}(-1,1)$
[…]
SNR $\rho$ $\Gamma(5,9)$
Table 1: Distribution of signal parameters used to generate the competition’s challenge. The Uniform distributions are defined by their lower and upper limits; the gamma distribution $(\Gamma)$ is expressed using the shape-scale parameterisation.
We describe the amplitude of a CW signal using the Signal-to-Noise ratio (SNR)
$\rho(h_{0},\psi,\cos\iota,\hat{n})=\sqrt{\frac{T_{\textrm{data}}}{S_{\textrm{n}}}}g(\psi,\cos\iota,\hat{n})h_{0}$ (3)
[…]
Blind CW searches usually quote their results in terms of population-based upper limits at a certain dismissal probability (usually 5% to 10%) on a constant-$h_{0}$ population [3, 5]. The false-alarm probability of a search is usually not computed. This complicates the comparison of Kaggle solutions to realistic searches on equal footing. Instead, we will base our comparisons on computing cost; concretely, we will assess the capability of Kaggle solution to discard broad parameter-space regions in such a way that signals detectable by a blind CW search are not discarded. To do so, it will suffice to re-scale the sensitivity estimates reported by the two reference searches to the dataset duration used in the competition.
The two reference searches we use in this challenge, based on the GCT [86] and Weave [87] pipelines, report their sensitivity estimates in terms of the average amplitude $\langle h_{0}\rangle$ at which 90% of an isotropically-oriented uniformly-sky-distributed signals would be detected. We relate this quantity to the corresponding averaged SNR as [96]
$\sqrt{\langle\rho^{2}\rangle_{\vec{n},\cos\iota,\psi}}=\frac{2}{5}\sqrt{\frac{T_{\textrm{data}}}{S_{\textrm{n}}}}\langle h_{0}\rangle\;.$ (4)
These pipelines operate by computing the semicoherent $\mathcal{F}$-statistic [20, 21]. Broadly, the dataset is split into segments with a duration $T_{\textrm{coh}}$ which are analysed independently using the $\mathcal{F}$-statistic, and then the results are combined to compute the semicoherent $\mathcal{F}$-statistic [23, 24, 22].
[…]
With this information, we choose to describe the SNR distribution using a Gamma distribution with shape parameter $k=5$ and scale parameter $\theta=9$, shown in Fig. 2. This distribution is such that about 20% of the signals are above the sensitivity of GCT and about 50% of the signals are above the sensitivity of Weave. About 6% of the signals display visible features ($\rho>80$).
[…]
GCT [86] 60 $1.3\times 10^{-25}$ 47 61
Weave [87] 240 $1.0\times 10^{-25}$ 36 47
Table 2: Summary of sensitivity estimates reported by the two searches considered in this work. These are computed by averaging the reported sensitivity estimates across the frequency band of the search.
Figure 2: Empirical cumulative distribution of $\rho\sim\Gamma(5,9)$. Vertical lines denote the re-scaled sensitivity of different search pipelines as discussed in the main text.
4 Competition results
The Kaggle competition was held from October 4th 2022 to January 4th 2023, and accumulated a total number of $\sim$30,000 submissions amongst 936 teams. Competitors were provided with a dataset as described in Sec. 3, and the task was to submit a ranking statistic for each sample in order to classify them as signal or noise.
Each submission was ranked according to the Area Under the Receiver Operating Characteristic (ROC) curve (AUC score). That is, the ranking statistics of a submission where used as thresholds $t$ to compute the false-dismissal probability (fraction of signal samples classified as noise, $p_{\mathrm{FD}}(t)$) at a given level of false-positive probability (fraction of noise samples classified as signals, $p_{\mathrm{FP}}(t)$). The AUC was then computed by integrating the area under the curve described by $(p_{\mathrm{FP}}(t),p_{\mathrm{FD}}(t))$. This ranking metric was the most suitable for our data-analysis problem given the available selection in the Kaggle platform.
The challenge actually makes use of two leaderboards, namely public leaderboard, which is visible throughout the duration of the competition, using a small fraction of the competition’s dataset to compute the AUC score, and the private leaderboard, which computes the AUC score using the complete dataset. This design allows participants to obtain a continuous assessment of their performance against other competitors, which is beneficial for a rapid iteration of the solution’s design, and minimises the chances of winning solutions ovefitting to a specific dataset. Additionally, the use of a random subset of the dataset minimises the probability of competitors successfully inferring the true classification of test data samples through repeated sumbmission [98, 99, 100].
We show in Fig. 3 the evolution of the leaderboard scores for the top 10 teams of the competition. Overall, one can distinguish two periods in the competition: From the start of the competition to about 10 days before the end the leaderboard was dominated by Space Coders. Ten days before the end, three more teams (JunKoda, PreferredWave, and BearWaves) managed to surpass their score. Overall, most teams converged rapidly into a sensible solution whose sensitivity remained broadly unchanged throughout the challenge.
Figure 3: Evolution of the top-10 teams throught the competition in terms the AUC score. The markers correspond to the private leaderboard score, while the shaded regions extend to the public leaderboard score. The cumulative number of submissions by these 10 teams is shown as a dotted line.
4.1 Top-scoring submissions
We divide the solutions proposed by the top-10 scoring teams into three broad classes: Track-Statistic, Track-Fitting, and Machine-Learning.
4.1.1 Track-Statistic
The first class contains solutions implementing a track-based detection statistic similar to those used in blind CW searches with short coherence times [26, 27, 101, 102, 28, 29]. Specifically, these methods assume a closed-form frequency evolution (i.e. a “frequency track” in the spectrogram) parameterised by a small number of parameters (such as frequency, spindown, and sky position) along which a detection statistic is accumulated. The search then consists in finding the parameters that maximise said detection statistic. This class includes the teams JunKoda [103] (1st), PreferredWave [104] (2nd), HiddenNeuralLayers [105] (5th), and Shun_PI [106] (6th). These methods present two main degrees of freedom, namely the detection statistic and strategy to cover the parameter space, which we summarize in Table 3. To address non-Gaussianities arising in real-data samples, all teams masked frequency bands in the spectrogram containing large deviations from Gaussian noise.
GPU Statistic Template placement
JunKoda (1st) Yes Maximum Dirichlet-weighed power [107] Square grid ($3.5\times 10^{8}$ templates)
PreferredWave (2nd) Yes Average power Uniform sampling ($3.3\times 10^{6}$ templates)
HiddenNeuralLayers (5th) No Maximum power Differential evolution (100 steps) [108]
Shun_PI (6th) No Maximum power Simulated-annealing [109].
Table 3: Design choices of the winning solutions presented by the teams in the Track-Statistic class of solutions.
The relative placement of these four teams in the leaderboard is consistent with their strategies. HiddenNeuralLayers and Shun_PL place templates in a data driven manner. While such a strategy has been proven effective in local analyses [7, 32, 33, 34], the breadth of the CW parameter-space makes them unsuitable for broad searches such as the ones here considered [7], hence their relatively lower placement. The difference between JunKoda and PreferredWave, on the other hand, can be ascribed to the significant difference in the number of evaluated templates ($10^{8}$ in JunKoda versus $10^{6}$ in PreferredWave) and the fact that JunKoda takes the maximum rather than the average power in the template bank [29].
4.1.2 Track-Finding
The second class introduces an extra degree of freedom with respect to Sec. 4.1.1, namely that the frequency evolution of the signal is not parameterised but inferred from the data. This approach is used by team SpaceCoders (4th), and bears some similarities to CW searches based on the Viterbi algorithm [53].
Specifically, the method uses dynamic programming to find a track maximising the accumulated SFT power. The track is constrained to contain a single inflection point, given the duration of the dataset and the expected signal morphology. Upon identifying a track, the signal’s parameters are estimated using a parametric fit in terms of frequency, spindown, and sky position. The optimisation is carried out using the Nelder-Mead algorithm [110], in a similar manner to the methods presented in Sec. 4.1.1.
The relative loss of SpaceCoders with respect to other track-based methods can be justified as follows: First, the number of tracks explored by the SpaceCoders search is exponentially greater than a closed-form track search [111, 53, 5]; this increases the expected background distribution [112] and, consequently, lowers the detection probability at a fixed amplitude. Second, the track partially relies on the signal displaying strong features within short timescales to clearly identify it. Due to the idiosyncrasy of CWs, this is only true for relatively strong signals.
4.1.3 Machine-Learning
The final class includes solutions using machine-learning strategies; that is, search pipelines in which some component is trained before producing a submission. This class comprises teams BearWaves [113] (3rd), KDL [^1] No available information about this team (7th), MCDigital [114] (8th), yu4u [115] (9th), and anonamename [116] (10th).
anonamename implemented a classifier based on a Convolutional Neural Network (CNN). Data samples are pre-processed following [116] and two outputs are produced: a detection statistic, and the estimated location of the signal in the spectrogram as a mask. To account for real data, line-like features were generated and included during training.
yu4u [115] implemented a similar strategy using a U-net architecture to estimate the location of the signal in the spectrogram and compute a detection statistic. As in the previous case, real data noise features were addressed by generating data containing line-like features during training.
MC Digital [114] combines two detection statistics using a CNN and the $\mathcal{F}$-statistic as computed by PyFstat [81]. Gaussian and real-data are addressed separately by training two different networks. The combination of CNN and $\mathcal{F}$-statistic results is done manually: samples with high $\mathcal{F}$-statistic values are classified according to the $\mathcal{F}$-statistic; samples with low $\mathcal{F}$-statistic values are classified according to the CNN.
Finally, BearWaves [113] presented a more involved approach. First, training data was simulated on-the-fly using PyFstat to avoid overfitting. To simplify the task of combining data from two detectors, data samples were time-averaged to obtain a common set of timsetamps. Data samples were analyzed by an ensemble of neural networks applying different processing strategies.
Based on their relative leaderboard positions, image-based machine-learning strategies tend to deliver less sensitive pipelines compared to track-based statistics unless a significant amount of resources is invested in data pre- and post-processing, as shown by BearWaves [113].
4.2 Reference submission: SOAP
We now describe the application of SOAP [53, 54, 55] to the competition’s dataset. SOAP is a fast, model-agnostic blind CW search capable of detecting a broad class of CW sources, even if they are affected by complicated physics such as spin-wandering, glitches, or timing noise [9]. This robustness, however, limits the sensitivity of SOAP to well-modelled isolated NSs, as they correspond to a minuscule portion of the probed parameter space [5].
Contrary to most blind CW searches, which identify interesting parameter-space regions where a signal may be present [32], SOAP’s main search stage assigns detection statistics to frequency bands by returning the most-significant track as constructed by the Viterbi algorithm [117]. This is consistent with the submission format expected in the Kaggle competition.
The agnosticity of SOAP makes it an ideal pipeline to conduct a first comparison between Kaggle solutions and real CW searches. More concretely, we require Kaggle solutions to return a greater AUC score than that achieved by SOAP, as otherwise their design choices would make them less performant than a search with minimal assumptions.
The SOAP analysis begins 1800s SFT power spectra, which are summed over one-day periods (48 SFTs). This daily summation process effectively average out antenna pattern modulation and increasing the SNR in each frequency bin. The search uses the line-aware statistic [53], a multi-detector approach that computes the Bayesian odds ratio between the signal hypothesis and the noise hypothesis including spectral lines [28]. This statistic specifically penalises instrumental line-like features in spectrograms, helping to distinguish genuine astrophysical signals from detector artifacts.
The line aware statistic incorporates three tunable parameters: the width of the prior distribution on signal SNR $\rho_{\rm{signal}}$, the width of the prior distribution on line SNR $\rho_{\rm{line}}$, and the ratio of prior probabilities between noise and line models $p(\mathrm{line})/p(\mathrm{noise})$. Following the search in LIGO O3 data [13], these parameters are set to $\rho_{\rm{signal}}=4.0$, $\rho_{\rm{line}}=10.0$, and $p(\mathrm{line})/p(\mathrm{noise})=0.4$.
Additional tunable parameters include the transition matrix probabilities that guide the algorithm’s frequency tracking behavior. Using the two-detector configuration, the transition matrix is configured to constrain the algorithm to search identical frequency bins across both detectors. This results in a three-element transition matrix defining probabilities for frequency transitions $(+T_{\textrm{SFT}}^{-1},0,-T_{\textrm{SFT}}^{-1})$ with equal values matching those used in the O3 search [13].
For each frequency band in the test dataset described in Section 3, a value of the line-aware statistic is produced for each band.
4.3 Comparison
The performance of the competition’s top-10 solutions is summarized using a ROC curve in Fig. 4, which shows the false-dismissal probability $p_{\mathrm{FD}}$ of a pipeline (or, equivalently, the detection probability $1-p_{\mathrm{FD}}$) as a function of false-positive probability $p_{\mathrm{FP}}$. The false-positive probability $p_{\mathrm{FP}}$ is defined as the fraction of noise samples that get classified as a signal; the false-dismissal probability $p_{\mathrm{FD}}$ corresponds to the fraction of signal samples that get classified as noise.
All solutions in the top 10 leaderboard produce an AUC above the value reported by the SOAP reference submission in Sec. 4.2, implying that all pipelines introduce cogent information into their assumptions when compared to an agnostic search.
The AUC score, however, aggregates results computed from the complete dataset across multiple false-positive probabilities. As a result, there is no single solution that dominates across all false-positive probabilities. Also, the fact that signals with different SNRs are combined together makes a direct comparison with standard CW search methods difficult.
Figure 4: ROC curves corresponding to the top 10 solutions of the competition. The legend lists the AUC scores corresponding to each solution. The thin black line corresponds to a non-discerning classifier with an AUC score of 0.5.
5 Comparison to CW searches
Blind CW searches operate by identifying interesting regions in a broad parameter-space to then follow them up with a more sensitive method or using a different dataset. The result of a search is a candidate (or collection of candidates) with a relatively-low associated significance that will be increased through follow-up steps if they are caused by CW signals. The solutions in this competition, due to its design, operate on data samples directly and return an associated false-positive probability $p_{\mathrm{FP}}$; that is, the significance they return is not associated to a specific template. Hence, these two approaches are not directly comparable.
On the other hand, the false-dismissal probability $p_{\mathrm{FD}}$, which is defined as the fraction of signals not detected by a method, is equivalent in both Kaggle solutions and CW searches, as in both cases it refers to the fraction of bands containing a signal that ended up labelled as noise.
This equivalence allows us to interpret Kaggle solutions as one-shot CW searches. Specifically, we consider the situation in which a Kaggle solution is configured to run a CW search at a relatively small $p_{\mathrm{FD}}$ so that the Kaggle solution discards a fraction of about $(1-p_{\mathrm{FP}})$ of the initial dataset, reducing the cost of the CW search, at a negligible sensitivity impact. The result of this first step would then be followed-up using other methods, depending on the intent of the search.
The cost of a standard CW search can be modelled simply as
$\mathcal{C}_{\mathrm{CW}}=c_{\mathrm{t}}N_{\mathrm{t}}$ (5)
where $c_{t}$ is the average computing cost per template and $N_{\mathrm{t}}$ is the number of templates evaluated by the search. $N_{\mathrm{t}}$ itself can be further expressed as
$N_{\mathrm{t}}=n_{\mathrm{t}}N_{\mathrm{bands}}$ (6)
where $n_{\mathrm{t}}$ is the average number of templates per band. The cost of a search including a Kaggle-based pre-processing step, on the other hand, can be expressed as
$\mathcal{C}_{\mathrm{K}+\mathrm{CW}}=c_{\mathrm{K}}N_{\mathrm{bands}}+p_{\mathrm{FP}}\mathcal{C}_{\mathrm{CW}}$ (7)
where $c_{\mathrm{K}}$ is the cost of running the model in a band and $p_{\mathrm{FP}}$ represents the fraction of bands flagged as containing a signal by the method. The relative cost ends up as
$\frac{\mathcal{C}_{\mathrm{K}+\mathrm{CW}}}{\mathcal{C}_{\mathrm{CW}}}=p_{\mathrm{FP}}+q_{\mathrm{K}}.$ (8)
where $q_{\mathrm{K}}=c_{\mathrm{K}}/(c_{\mathrm{t}}n_{\mathrm{t}})$. As a result, Kaggle solutions will provide an advantage if they are computationally cheaper ($q_{\mathrm{K}}<1$) and can reach a sufficiently low $p_{\mathrm{FP}}$ at a low $p_{\mathrm{FD}}$ level.
Strictly speaking, $c_{t}$ and $n_{t}$ depend on the specific implementation of a CW search [6, 118, 24, 28, 25, 29]. We here instead take an educated guess based on the latest result in the literature. The latest GCT searches conducted in LIGO O2 and O3 data [86, 12] use $n_{\mathrm{t}}\sim 10^{12}-10^{15}$, depending on the frequency band. Similarly, the computing cost of the $\mathcal{F}$-statistic is about $c_{\mathrm{{t}}}\sim(10-100)\,\mathrm{ns}$ [22, 86, 12, 87]. Thus, $c_{\mathrm{t}}n_{\mathrm{t}}\sim(10^{4}-10^{7})\,\mathrm{s}$.
[…]
We finally evaluate the performance of each Kaggle solution in restricted SNR ranges matching the two reference searches previously discussed, namely $\rho\in[40,50]$ corresponding to Weave and $\rho\in[50,60]$ corresponding to GCT. The results are shown in Fig. 5 for Gaussian and real-data samples. The number of signal samples within said ranges limits the resolution of false dismissal probabilities to about $(1-2)\%$, and consequently we evaluate $p_{\mathrm{FP}}$ at $1-p_{\mathrm{FD}}=0.98$. Similarly, the precision on $p_{\mathrm{FP}}$ is about $10^{-3}$; given the computed value of $q_{\mathrm{K}}$, this implies the computational gains are dominated by $p_{\mathrm{FP}}^{-1}$.
Figure 5: False dismissal probability as a function of false-positive probability for a restricted range of SNR. The upper row shows the results using Gaussian noise samples, while the lower row uses only real-data samples. The shaded regions show binominal errorbars.
For the Weave range ($\rho\in[40,50]$), in the case of Gaussian noise, JunKoda reaches a false dismissal probability $p_{\mathrm{FD}}=0.02$ at $p_{\mathrm{FP}}\sim(10^{-2}-10^{-1})$, while Preferred Wave performs at $p_{\mathrm{FP}}\sim(0.2-0.4)$. All other solutions achieved the allowed false dismissal probability at $p_{\mathrm{FP}}\gtrsim 0.5$. The potential gain in this case is from a factor of about 5 (Preferred Wave) to one to two orders of magnitude (JunKoda). The situation in the case of real noise remains comparable within the uncertainty.
For the GCT range ($\rho\in[50,60]$), in Gaussian noise, JunKoda, Preferred Wave and Space Coders reach $p_{\mathrm{FP}}\sim 10^{-3}$ at the specified false dismissal probability level. The next relevant team is yu4yu (9th), at $p_{\mathrm{FP}}\sim 3\times 10^{-2}$. The remaining solutions about $p_{\mathrm{FP}}\sim 10^{-1}$.
Despite the increased uncertainty, the situation in real-data samples changes slightly: JunKoda stays at $p_{\mathrm{FP}}\sim 10^{-3}$; Preferred Wave worsens up to $p_{\mathrm{FP}}\sim 7\times 10^{-3}$, and Space Coders loses an order of magnitude up to $p_{\mathrm{FP}}\sim 3\times 10^{-2}$. yu4u maintains a comparable performance to the Gaussian case. The relative loss of PreferredWave with respect to JunKoda can be ascribed to the use of average power in a template bank: In non-Gaussian data, the average template-bank power tends to rise due to the presence of extreme values in the background, even if most of the problematic bins have been removed. Similarly, the loss of Space Coders can be related to the fact that their method builds-up a frequency-evolution track in a data-driven manner, an thus can get biased due to local maxima in the spectrogram. Finally, we highlight the performance of Hidden Neural Layers (5th), which reaches $p_{\mathrm{FP}}\sim 10^{-3}$ surpasing all other teams except JunKoda. This is likely a consequence of their approach to isolate and remove potential glitches [105].
6 Discussion
CW signals are one of the potential next discoveries of gravitational-wave astronomy. Currently their detection is complicated by our ignorance on the specific emission mechanism at play, which gives rise to a broad parameter-space that requires significant computing budgets to be effectively covered. Fundamentally, the sensitivity of CW searches for unknown sources is limited by the available computing [6, 24, 5].
In this work, we explored the use of data-analysis competitions to discover new search strategies. Specifically, we framed CW searches as a classification problem and deployed a Kaggle competition to find the most effective strategy to identify data samples harbouring simulated CW signals. Kaggle [73] is a online data-analysis competition platform powered by Google hosting a broad community of data-analysis practitioners with a wide variety of backgrounds. By tapping into a new, broad community of experts, we intended to asses the performance of data-analysis methods from other field when applied to the problem of detecting CW signals. This approach is similar to mock data challenges previously developed for other gravitational-wave use cases [57, 58, 60, 59].
The challenge lasted for 3 months and attracted more than 1,000 participants. The specific task was to classify 8,000 data samples (containing Gaussian noise, real detector data, and possibly simulated CW signals). Each submission was ranked according to the Area under the ROC curve metric. Upon completion, we contacted the top 10 solutions to understand their approach and implementation.
Overall, the competition was dominated by pipelines based on semicoherent matched filtering, in a similar fashion to [26, 27]. The proposed solutions, however, introduced several improvements with respect to well established methods.
First, the evaluation of detection statistics was vectorised over multiple waveforms using GPUs without using any signal approximation. This contrasts with most GW GPU-accelerated methods so far, which either parallelize the evaluation of a single template [22, 119] or compute an approximated version of the detection statistic [28, 120]. The potential improvement in computational cost and complexity of search pipelines inspired the development of fasttracks [29], a unified, GPU-accelerated engine to compute CW detection statistics using SFTs for general CW waveform models.
Second, the winning solution [103] made use of alternative filtering strategies, similar to those previously outlined in Ref. [107]. While these specific statistics have not been applied to short-coherence CW searches, their efficiency allowed us to extend the use-case of efficient SFT-based detection statistics to the search for binary neutron stars and massive black-hole binaries [121] in next-generation detectors (Einstein Telescope [122], Cosmic Explorer [123], LISA [124]).
Solutions accounting for the presence of non-Gaussian data did so in a relatively simple manner, either by masking problematic regions in the spectrogram [103, 104, 105, 106] or by generating non-Gaussian training data [113, 114, 115, 116]. This is comparable to the practical approach taken in CW searches, where non-Gaussianities can be identified as sudden spikes in the accumulated power [28, 29].
Our analysis suggests that, for datasets compared to those used in this competition, the use of Kaggle solutions could reduce the number of frequency bands to be evaluated by one to two orders of magnitude.
The relevance of approaches based on alternative techniques, such as machine learning, becomes apparent when analyzing specific subsets of SNR. For example, for moderately strong signals $\rho\in[50,60]$, we find that U-net architectures [115] are able to compete with track-based statistics and show a negligible performance degradation when applied to non-Gaussian data.
Despite the limited duration of this challenge, the obtained solutions have the potential to affect future CW searches not only by improving the efficiency and deployment of track-based searches [29, 121], but also by spurring approaches prescinding from template banks all together. This latter point is particularly interesting, as the computing cost of a CW search is dominated by the evaluation of statistics on a template bank to numerically marginalise with respect to the signal parameters to compute a Bayes factor [76, 77, 78]; this Kaggle competition, on the other hand, allows for pipelines operating directly on data samples, so that the method itself constitutes an approximation of the marginalised Bayes factor. The solutions here collected, together with those proposed from within the community [48, 125, 54, 49, 55, 52, 50], bear witness of the potential of alternative approaches to revolutionize the search for blind CW sources.
Although the dataset is shorter than typical production searches, the chosen duration (4 months) remains a valid regime for candidate generation. Given the expected cadence of forthcoming LVK observing runs, these results support a strategy based on analysing each run shallowly to produce candidate lists, with significance subsequently accumulated across runs—akin to the “fresh-data mode” discussed in Ref. [126]. Such implementation will be presented elsewhere.
This Kaggle competition is the first public, large-scale exposure of the CW detection problem to a wide community of data-analysis practitioners, and a logical step beyond established closed-door mock-data challenges [57, 58, 60, 59] to foster cross-pollination across fields. Its impact is already evident in new GPU-accelerated methods [29] and in applying CW-inspired detection statistics to other domains [121]. The materials developed for this competition, including tutorials for generating and pre-processing CW data with PyFstat [81], and the dataset itself [61], are publicly available to support further development of alternative methods aimed at the first CW detection.
Acknowledgements
We thank Rafel Jaume, David Keitel, Narenraju Nagarajan, Alicia M. Sintes, and Daniel Williams for support and discussions, and Christopher Berry, Lorenzo Mirasola, and Karl Wette for comments on the manuscript. We thank Kaggle/Google for contributing to the competition prize, and acknowledge support from COST Action CA17137 (G2Net). RT thanks Graham Woan and the Institute for Gravitational Research at the University of Glasgow for their hospitality during the development of this work. RT is supported by ERC Starting Grant No. 945155–GWmining; Cariplo Foundation Grant No. 2021-0555; MUR PRIN Grant No. 2022-Z9X4XS; MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ); Italian-French University (UIF/UFI) Grant No. 2025-C3-386; the ICSC National Research Centre funded by NextGenerationEU; the Universitat de les Illes Balears (UIB); the Spanish Agencia Estatal de Investigación grants PID2022-138626NB-I00, RED2024-153978-E, RED2024-153735-E, funded by MICIU/AEI/10.13039/501100011033 and the ERDF/EU; and the Comunitat Autònoma de les Illes Balears through the Conselleria d’Educació i Universitats with funds from the European Union - NextGenerationEU/PRTR-C17.I1 (SINCO2022/6719) and from the European Union - European Regional Development Fund (ERDF) (SINCO2022/18146). MJW is supported by the Science and Technology Facilities Council [2285031, ST/X002225/1, ST/Y004876/1], and the University of Portsmouth. JB and CM are supported by STFC grant ST/V005634/1. Computational work was performed at MareNostrum5 with technical support provided by the Barcelona Supercomputing Center (RES-FI-2024-3-0013, RES-FI-2025-1-0022), and at CINECA with allocations through INFN and Bicocca. This research has made use of data or software obtained from the Gravitational Wave Open Science Center (gw-openscience.org), a service of LIGO Laboratory, the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This paper has been assigned document number LIGO-P2200295.
Appendix A First Kaggle Competition
The G2Net Kaggle challenge (G2Net Gravitational Wave Detection), launched on 30 June 2021, marked the first public data competition in gravitational-wave astrophysics aimed specifically at the machine learning community. Its goal was to explore the capabilities of machine learning classifiers for detecting transient GW signals from stellar-mass binary black hole (BBH) mergers. Competitors were tasked with predicting the presence or absence of a BBH signal within short time-series recordings of simulated detector data. Similarly to the CW challenge discussed in this paper, the performance was also evaluated using the AUC, computed from participants’ probability estimates of signal presence in the test set.
The dataset comprised simulated measurements from the three-detector network of advanced interferometers: LIGO-Hanford, LIGO-Livingston, and Virgo. Each training example consisted of three time series, corresponding to the three detectors, covering two seconds sampled at 2048 Hz. Some samples contained only Gaussian noise, while others were augmented with simulated GW signals. The simulated BBH signals were generated using IMRPhenomPv2 [127] and varied across 15 astrophysically relevant parameters, including component masses, spins, binary orientation, sky location, distance, polarization, merger phase, and arrival time. Although the distributions of these parameters reflected realistic astrophysical populations, the overall signal amplitudes were adjusted to produce a mix of challenging and easier detection cases. Most signals were designed to lie near the typical sensitivity threshold of matched-filter searches, around an SNR of 8.
The challenge attracted broad participation, with 1,501 individuals forming 1,219 teams. Competitors applied a variety of machine learning approaches, from convolutional and recurrent neural networks to hybrid architectures, often incorporating data augmentation and ensemble methods to improve generalization. The competition validated claims in the literature [36, 128, 129] that deep learning methods could achieve competitive detection performance on realistically simulated multi-detector GW datasets. Moreover, the challenge provided a benchmark for future efforts in applying modern machine learning techniques to gravitational-wave detection, highlighting both the potential and limitations of data-driven approaches in this domain.
Figure 6: Evolution of the top-3 teams throughout the competition in terms the AUC score. The markers correspond to the private leaderboard score, while the shaded regions extend to the public leaderboard score. The evolution of the private leaderboard score for other teams is shown in grey.
Figure 7: Left: ROC curves corresponding to the top-3 solutions of the first Kaggle competition. The legend lists the AUC scores corresponding to each solution. The black line corresponds to a non-discerning classifier with an AUC score of 0.5. Right: Efficiency as a function of network matched SNR (computed using the injection parameters) at a FAR of 1 per day for the top-3 solutions of the first Kaggle competition. We use 50 bins with logarithmic spacing between SNR 1 and 100 to produce the curve. The black line shows the cumulative fraction of signals in the original dataset as a function of SNR.
In Fig. 6, we show the evolution of the AUC scores achieved by all participating teams over the course of the competition. Most final solutions clustered tightly around an AUC of $\sim 0.88$. The highlighted curves for the top three solutions indicate that their initial approaches could readily achieve AUCs of $\sim 0.875$. The ROC curves in Fig. 7 provide additional insight into the relative performance of the top three solutions. Of particular note is that the winners, KDL, were able to gain an advantage over 2nd place EventHorizon and 3rd place GOGOGO, achieving marginally higher overall probability of detection but, importantly, substantially better performance at the lowest false-positive rates. Figure 7 also provides the detection efficiencies of each of the top three teams as a function of network SNR at a fixed false positive probability equivalent to 1 per day. It is interesting to see clear distinction in sensitivity between the teams at efficiencies $\sim 0.5$ but the KDL team achieve efficiencies of 1 at significantly lower SNR. We also note that the 2nd place efficiency curve indicates a potential dataset bias or extrapolation failure where under-representation in training can lead to extremely loud signals being outside the learned feature distribution and hence classified incorrectly as noise.
The top three teams share several key strategies that contributed to their performance. They emphasised careful preprocessing of the waveform, including whitening, bandpass filtering, and mitigation of edge effects, while selectively using augmentation techniques such as Gaussian noise, waveform flipping, and time shifts to improve generalisation. Each team used 1D convolutional architectures as primary frontends, often in combination with 2D CNN backends or hybrid models, and recognised the importance of channel-specific processing—sharing weights for similar detectors while handling others separately. Synthetic data generation and pre-training played a critical role in avoiding overfitting and providing access to auxiliary targets, and all groups used ensembles of multiple models to further boost robustness. Finally, iterative approaches like pseudo-labeling, test-time augmentation, and progressive refinement were widely employed to exploit unlabelled data and maximize leaderboard performance.
The winning team (KDL) overcame the main competition challenge — overfitting— by generating large amounts of realistic synthetic data for pre-training, which also allowed the use of hidden parameters as auxiliary targets. Their model design included learnable frontends for each channel, transformation of 1D data into richer time-frequency or feature spaces, and lean encoders to limit overfitting. Key innovations included a custom 1D CNN with large kernels, recognition that 1D models outperform 2D CNNs, and extensive use of synthetic data for pre-training. They also observed fundamental limitations: that about 30% of true signals are inherently undetectable, some noise samples are always misclassified as signals, and training on difficult samples can invert model predictions.